OpenStudy (anonymous):

WHOEVER ANSWERS THIS GETS A MEDAL AND A FAN <3 Ok so they're asking whether these equations are a direct variation, and if they are, whats the constant variation? 4y=x & y/x=3

OpenStudy (campbell_st):

so a direct variation equation is in the form y = kx k is the constant of variation... so in each equation, make y the subject

OpenStudy (campbell_st):

so for the 1st equation, divide both sides by 4 in the 2nd equation. multiply both sides of the equation by x... hope it helps

OpenStudy (anonymous):

So basically I just flip them around? For the first one would K = 4 then..?

OpenStudy (campbell_st):

not quite 4y = x divide both sides by 4 to get y = ?x what would ? be..? it's not 4

OpenStudy (anonymous):

Hmm....this confuses me but I suppose then it's just not a direct variation then?

OpenStudy (campbell_st):

well its \[\frac{4y}{4} = \frac{x}{4} \] which simplifies to \[y = \frac{1}{4} x\] so what is the constant of variation

OpenStudy (anonymous):

OH! so then the constant of variation is 1/4...right..?

OpenStudy (anonymous):

How about the second equation? y/x=3 Would it look something like this..? y=3*x

OpenStudy (unklerhaukus):

yep

OpenStudy (campbell_st):

correct for both questions

OpenStudy (anonymous):

thanks!